Using your Calculator - calculate the bond equivilent yield of this bond (quoted rate):
Answer should be in 3 decimal places and whole numbers: ex: 6.125 not 6.125% or 0.06125

Price $89.5

Coupon: 5.5%

Maturity Date: 9/15/2026

Today's Date: 9/15/2018

Compounding/Payments: Semi-Annual
Answer #1

Using financial calculator
PV=-89.5
FV=100
PMT=5.5%*100/2=2.75
N=(2026-2018)*2=16
CPT I/Y=3.625%

Hence, bond equivalent yield=3.625%*2=7.25%